E = I(R + r) .... (i)

V = IR ...(ii)

Use equations (i) and (ii) to solve for r by eliminating I and write this equation the form:   y = mx + b

Identify the independent variable and the dependent variable
Answer #1

Here E=I (R47) V=IR (0) C) (175) Forom equation (ii) we get I= ☆ Put the valve at I jn equation (i) we get E= Ř (R+r) = E= VR